OpenStudy (anonymous):

Find domain of the function 1/(x-|x|)

OpenStudy (anonymous):

\[\frac{ 1 }{ x-|x|}\]

OpenStudy (anonymous):

x \[x \in \left\{ r \right\}-\left\{ -1 \right\}\]

OpenStudy (anonymous):

What is r?

OpenStudy (anonymous):

@manojbalaji1

OpenStudy (anonymous):

sorry it is R not r where r is set of real numbers

OpenStudy (anonymous):

@Palmo4ka

OpenStudy (anonymous):

okay, i understood. Thank you! Can you help me with else one question? \[y=\frac{ 1 }{ x+|x| }\] @manojbalaji1